# Non-strict Singularity of Optimal Sobolev Embeddings

### v1

@@ 4. Optimal embeddings of “nonlimiting” $\Lambda$ spaces and their non-strict singularity
thanks to (4.31) and (4.32). Furthermore, thanks to the absolute continuity of $\|\cdot\|_{\Lambda^{q}_{w_{opt}}(\Omega)}$, the monotone convergence theorem, (4.36), and the absolute continuity of the integral, we can find $40<\tilde{a}<a_{1}/$ 4 and $\tilde{B}_{2}=B(x_{0},\tilde{a}^{1/n}R)$ so small that
@@
and that the inequalities (4.37)–(4.40) still hold with $\tilde{B}_{2}$ and $\tilde{a}$ replaced by $B_{2}$ and $a_{2}$, respectively. Having found $a_{2}$ and $B_{2}$, we now use Theorem 3.6 again to find functions $f_{2}=c_{2}\chi_{(a_{2}/2,a_{2})}\in\mathscr{S}_{a_{2}/2,a_{2}}$ for some $c_{2}>0$ and $u_{2}=u_{f_{2}}\in V^{m}_{0}X(\Omega)\cap\mathcal{C}^{m-1}(\Omega\setminus\{x_{0}\})$ such that (4.32)–(4.35) hold with $f_{1}$, $u_{1}$, and $B_{1}$ replaced by $f_{2}$, $u_{2}$, and $B_{2}$, respectively.
@@
(4.44) the support of $u_{j}$ is the closed ball $\text{the closed ball $\bar{B}_{j}$},$
@@
(4.51) $(1+\varepsilon_{1})\|u_{1}\chi_{B_{j}\setminus B_{j+1}}\|_{\Lambda^{q}_{w_{opt}}(\Omega)}$ $>C_{2},$
@@
[…]

### v2

@@ 4. Optimal embeddings of “nonlimiting” $\Lambda$ spaces and their non-strict singularity
thanks to (4.31) and (4.32). Furthermore, thanks to the absolute continuity of $\|\cdot\|_{\Lambda^{q}_{w_{opt}}(\Omega)}$, the monotone convergence theorem, (4.36), and the absolute continuity of the integral, we can find $0<\tilde{a}<a_{1}/4$ and $\tilde{B}_{2}=B(x_{0},\tilde{a}^{1/n}R)$ so small that
@@
and that the inequalities (4.37)–(4.40) still hold with $\tilde{B}_{2}$ and $\tilde{a}$ replaced by $B_{2}$ and $a_{2}$, respectively. Having found $a_{2}$ and $B_{2}$, we now use Theorem˜3.7 again to find functions $f_{2}=c_{2}\chi_{(a_{2}/2,a_{2})}\in\mathscr{S}_{a_{2}/2,a_{2}}$ for some $c_{2}>0$ and $u_{2}=u_{f_{2}}\in V^{m}_{0}X(\Omega)\cap\mathcal{C}^{m-1}(\Omega\setminus\{x_{0}\})$ such that (4.32)–(4.35) hold with $f_{1}$, $u_{1}$, and $B_{1}$ replaced by $f_{2}$, $u_{2}$, and $B_{2}$, respectively.
@@
(4.44) the support of $u_{j}$ is $\text{the closed ball $\bar{B}_{j}$},$
@@
(4.51) $(1+\varepsilon_{1})\|u_{j}\chi_{B_{j}\setminus B_{j+1}}\|_{\Lambda^{q}_{w_{opt}}(\Omega)}$ $>C_{2},$
@@
$\begin{aligned} \leq 2^{q-1}\sum_{k=1}^{N}\Bigg(&|\alpha_{k}|^{q}|c_{k}|^{q}\int_{2a_{k+1}}^{2a_{k}}\chi_{(0,a_{k}/2)}(t/4)w(t)\mathop{}\!\mathrm{d}{t}\\ &+\int_{2a_{k+1}}^{2a_{k}}\Big(\sum_{j=1}^{k-1}f_{j}\Big)^{*}(t/4)^{q}w(t)\mathop{}\!\mathrm{d}{t}\Bigg)\end{aligned}$
(4.60) $\begin{aligned} \leq 2^{q-1}\sum_{k=1}^{N}\Bigg(&|\alpha_{k}|^{q}c_{k}^{q}W(2a_{k})\\ +&\int_{2a_{k+1}}^{2a_{k}}\Big(\sum_{j=1}^{k-1}f_{j}\Big)^{*}(t/4)^{q}w(t)\mathop{}\!\mathrm{d}{t}\Bigg).\end{aligned}$
@@
Acknowledgment
The authors would like to thank the referee for carefully reading the paper and their valuable comments.